OpenStudy (anonymous):

secant(-30) as fraction

OpenStudy (anonymous):

-30 = -pi/6, and secant is 1/cos(-30), and cos (-pi/6)= -pi/6=-sqrt(3)/2), so 1/ -(sqrt(3)/3, so -3/sqrt(3)

OpenStudy (callisto):

Not really correct... secant(-30) = sec (30) since cos (-30) = cos 30 sec30 = 1/ (cos30) = 1/ (√3 /2) = 2/√3

OpenStudy (callisto):

Rationalise the answer if you need